OpenStudy (anonymous):

Sammy's mother bought 2 1/2 pounds of blueberries on monday. Sammy ate 1/4 of the blueberries before be went to bed. How many pounds did he eat on monday? 2. How many pounds were left? 3 on tuesday sammy ate 1/3 of the remaining blueberries how may pounds did he eat on tuesday? 4 did sammy eat more on monday or tuesday explain your reasoning. Thanks!

OpenStudy (anonymous):

2.5*0.25 = 5/8 pound 2.5 - .625 = 1 7/8 pounds 1 7/8 * 1/3 = 5/8 pound He ate the same amount